# Measuring Psychological Depth in Language Models

### v1

@@ 9 Limitations and Risks
Sourcing Human-written Stories from Reddit. Using human-written content from Reddit’s r/WritingPrompts has potential limitations. First, we cannot guarantee that all selected stories are fully written by humans. Despite community rules explicitly prohibiting AI-generated content, our study shows that many publicly available LLMs can sufficiently mimic human creative writing to evade detection of the time on average. Second, while Reddit’s voting system and large user base provide a reasonable signal of writing quality, it may not represent the highest caliber of human writing. To address this, we introduced writing quality levels to facilitate comparisons with some of the platform’s best content. Future research could further this by identifying more reliable sources of high-quality stories.
@@
Prompt Engineering. Engineering effective prompts is an active area of research Zhou et al. (2023b); Pryzant et al. (2023) and the templates we carefully reviewed and iteratively improved are still likely to be sub-optimal. We will open-source the full pipelines used for both story generation and evaluation. We encourage future work to refine these prompts by incorporating potentially compatible techniques Wei et al. (2023); Liu et al. (2023); Li et al. (2023) and current best practices Bsharat et al. (2024).
@@ A.2 Prompt Premises
As an additional precaution against potential plagiarism, we ensured that the selected premises were posted after the reported training data cut-off dates for GPT-4 (September 2021) OpenAI et al. (2023) and Llama-2 (September 2022) Touvron et al. (2023) to maximize the likelihood of generating genuinely new stories.
@@ A.5 Study Details
This study was reviewed by an IRB and determined to be "exempt" due to the nature of human involvement — i.e. rather than collecting information about the participants themselves, their involvement focused on story annotations. Consent was provided by continuing with the study after our tutorial and instructions relayed that anonymized annotations may be used to facilitate validation of our results and future work.

### v2

@@ 9 Limitations and Risks
Sourcing Stories from Reddit. Using human-written content from Reddit’s r/WritingPrompts has potential limitations. First, we cannot guarantee that all selected stories are fully written by humans. Despite community rules explicitly prohibiting AI-generated content, our study shows that many publicly available LLMs can sufficiently mimic human creative writing to evade detection of the time on average. Second, while Reddit’s voting system and large user base provide a reasonable signal of writing quality, it may not represent the highest caliber of human writing. To address this, we introduced writing quality levels to facilitate comparisons with some of the platform’s best content. Future research could further this by identifying more reliable sources of high-quality stories.
@@
Prompt Engineering. Engineering effective prompts is an active area of research Zhou et al. (2023b); Pryzant et al. (2023) and the templates we carefully reviewed and iteratively improved are still likely to be sub-optimal. We have open-sourced the full pipelines used for both story generation and evaluation. We encourage future work to refine these prompts by incorporating potentially compatible techniques Wei et al. (2023); Liu et al. (2023); Li et al. (2023) and current best practices Bsharat et al. (2024).
@@
10 Acknowledgements
This research is partly supported by a National Science Foundation CAREER award #2339766, a Meta/FAIR-sponsored research award, a Google Research Scholar grant, a Simons Investigator Award, NTT Research, NSF Grant #2333935, and the Symantec Chair of Computer Science. We are grateful to the reviewers and to people like Aaron Hatrick, Darrin Murray, and Sean Gildersleeve for their valuable feedback and insightful discussions, which helped shape the development of the ideas presented in this manuscript.
@@ A.1 Dataset Quality Control Measures
Ascertaining Human Authorship on Reddit. To better understand the likelihood of AI vs human authorship, we took a sample of 134 stories and passed them into Ghostbuster Verma et al. (2024), a tool with a reported 99% F1 accuracy in detecting AI-generated content. The results shown in Table 5 suggest that our human stories have a very low probability of AI authorship (10% average). While many of the LLMs are reliably detected, GPT-4 is noticeably challenging to detect (only 3.5% more likely to be AI than the human average). These results are consistent with our human study where our annotators perceived GPT-4 to be only slightly less human-like than the stories sourced from Reddit. Therefore, we still cannot definitively rule out the possibility of Redditors using GPT-4 to author their stories in part or full. The same is likely true for any recent writing contest, where monetary prizes may incentivize AI use. Perhaps the only way to guarantee sole human authorship is to directly observe study participants writing unassisted. However, this would complicate the experimental setup significantly, add more time, make it more difficult to attain high-quality writing, and estimate that quality as well.
| Author | # of Stories | Average AI Probability | Std Deviation |
| Human-Advanced | 15 | 0.1167 | 0.0927 |
| Human-Intermediate | 15 | 0.0907 | 0.0493 |
| Human-Novice | 15 | 0.0967 | 0.0613 |
| Human (Average) | 45 | 0.1013 | 0.0695 |
| Llama-2-7B | 19 | 0.7637 | 0.2299 |
| Llama-2-13B | 18 | 0.8744 | 0.1865 |
| Vicuna-33B | 16 | 0.8000 | 0.2273 |
| Llama-2-70B | 18 | 0.8956 | 0.0733 |
| GPT-4 | 18 | 0.1356 | 0.1917 |
| LLM (Average) | 89 | 0.6922 | 0.3411 |
These findings suggest that traditional indicators of writing quality are becoming increasingly unreliable for distinguishing human authorship from AI-generated content. Advanced models like GPT-4 have demonstrated a high level of proficiency in generating fluent, coherent, and contextually appropriate text across a wide range of topics. Consequently, recent research in AI text detection has shifted towards identifying discriminative textual-linguistic features, such as those explored by approaches like Ghostbuster Verma et al. (2024). We view such efforts as a form of stylometric analysis, akin to authorship attribution studies Neal et al. (2017), which aim to assess the likelihood that a document was authored by a specific individual based on stylistic traits. When reduced to the binary classification of "human" versus "AI" authorship, this task simplifies the number of author options but complicates the feature analysis due to the considerable variability within both categories.
Future research could benefit from a deeper exploration of established stylometric techniques and the development of experimental setups tailored to AI-generated content. Given the wide variation in LLM capabilities, it may be more effective to analyze their stylistic feature distributions separately. For instance, we observe that syntactic patterns such as adverbial present-participle phrases modifying a main clause (e.g., "We aimed to teach students effectively, closing the achievement gap.") appear frequently in GPT-4’s outputs. While the predictive power of individual features may be limited, combining multiple indicators could result in a robust classifier. Ultimately, a more nuanced understanding of these and other linguistic features will enhance AI detection methodologies and help ensure authorship credit where it is important to differentiate.
@@ A.2 Prompt Premises
As an additional precaution against potential plagiarism, we ensured that the selected premises were posted after the reported training data cut-off dates for GPT-4 (September 2021) Achiam et al. (2023) and Llama-2 (September 2022) Touvron et al. (2023) to maximize the likelihood of generating genuinely new stories.
@@ A.5 Study Details
This study was reviewed by an IRB and determined to be exempt due to the nature of human involvement — i.e. rather than collecting information about the participants themselves, their involvement focused on story annotations. Consent was provided by continuing with the study after our tutorial and instructions relayed that anonymized annotations may be used to facilitate the validation of our results and future work.
@@
Controlling for Annotator Fatigue. The stories were organized into batches of 20 to offset the possible impact of annotator fatigue. We anticipated annotator fatigue based on our internal annotation efforts early in the project and decided to encourage (not require) breaks between each batch. In our exit survey, we asked the following questions:
“Did you at any point feel fatigued over the course of annotating the short stories? (Please answer honestly. Study rewards are not affected by this question)”
Responses:
4 Yes
1 Maybe
“If you answered yes, do you think you would have provided different annotations if you were not fatigued? (Please answer honestly. Study rewards are not affected by this question)”
Responses:
4 No
1 Maybe
Overall, the majority of participants (4 out of 5) reported feeling fatigued during the annotation process, but none believed it significantly affected their annotations, with only one expressing some uncertainty.